Grade 10Physical Chemistry

if a solid compound AB having structure of rock salt has A ions in ccp arrangement and one of the diagonals of the cube is removed....what is the resultant formula

Rex Xii
8 Years agoGrade 10

1 Answer

Shivanjayraj
8 Years ago
Contribution of A after removal =4- (2×1/8)=15/4
Contribution of B after removal =4-1=3
A:B=15/4×3=5:4
A5B4